OpenStudy (anonymous):

The endpoints of segment EF are E(-2, 3) and F(5, -3). Find the coordinates of the midpoint of EF.

OpenStudy (lgbasallote):

\[\huge \left(\frac{x_1 + x_2}{2} \; , \; \frac{y_1 + y_2}{2} \right)\] \[\huge \implies \left(\frac{-2 + 5}{2} \; , \; \frac{3 - 3}{2} \right)\] does that help?

OpenStudy (anonymous):

yes and then what do i do next

OpenStudy (lgbasallote):

simplify that

OpenStudy (anonymous):

i got 3/2 0/2

OpenStudy (lgbasallote):

that's right (3/2 , 0)

OpenStudy (anonymous):

(3, 0) (½, 2) (0, 5) (1 ½, 0) these are the answers

OpenStudy (lgbasallote):

3/2 = 1 1/2

OpenStudy (lgbasallote):

if you simplify it

OpenStudy (anonymous):

so the answer is 1 1/2 and 2

OpenStudy (anonymous):

or 1 1/2 and 0

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

what is the y-coordinate in the one i wrote a while ago? 2 or 0?

OpenStudy (anonymous):

3/2, 0